OpenStudy (anonymous):

what equation that represents the line passing through the point (-5, 1) with a slope of 3/2.

OpenStudy (jackellyn):

y-y1=m(x-x1) use this formula to find the equation. m=slope

OpenStudy (anonymous):

Explanation✐ Use the slope formula.... y - y1 = m(x - x1) where... (x,y) and (x1,y1) are points on the line m is the slope Let... (2,1) = (x1,y1) m = 3 Substitute those to the formula and solve for y. y - 1 = 3(x - 2) y - 1 = 3x - 6 You get: y = 3x - 5 I hope that helps!

OpenStudy (anonymous):

use M=slope to find the formula.(:

OpenStudy (anonymous):

okay :) thank you all!

OpenStudy (anonymous):

welcome(: